OpenStudy (anonymous):

Elbert purchased a living room set for $5,021 using a 12-month deferred payment plan. The interest rate after the introductory period is 15.60%. A down payment of $375 is required as well as a minimum monthly payment of $122. What is the balance after the introductory period if only the minimum payment is made until then?

OpenStudy (anonymous):

Can someone help me please........

OpenStudy (anonymous):

do you have a formula for this? how long is the "introductory period"?

OpenStudy (anonymous):

\(375+12\times 122=1839\) so if the introductory period (i.e. no interest has been applied) is 12 months, the total remaining is \(5021-1839\)

OpenStudy (anonymous):

its introductory period is 15.60%.

OpenStudy (anonymous):

is that what you mean sorry

OpenStudy (anonymous):

well the choices i have to pick from are $4,023.78 $3,960.91 $5,302.91 $5,424.91

OpenStudy (anonymous):

i don't know how to do this other than step by step

OpenStudy (anonymous):

is ok dont worry about=D your awsome for trying to help either way

OpenStudy (anonymous):

if you make the down payment, you have 4646 remaining total after one month will be \(4646(1+\frac{.156}{12})=4706.40\) then you pay 122 have a balance of \(4706.40-122=4584.40\)

OpenStudy (anonymous):

now repeat for month two

OpenStudy (anonymous):

after month two you will owe 4522 http://www.wolframalpha.com/input/?i=4584.4%281%2B.156%2F12%29-122

OpenStudy (anonymous):

it is decreasing slowly so answers 3 and 4 are wrong i think

OpenStudy (anonymous):

i got 4643=D

OpenStudy (anonymous):

if you have a formula for this you should use it

OpenStudy (anonymous):

oh i didnt minus

OpenStudy (anonymous):

no the first number is \(5021-375=4646\) for sure

OpenStudy (anonymous):

this is what i get after month 3 http://www.wolframalpha.com/input/?i=4522%281%2B.156%2F12%29-122

OpenStudy (anonymous):

to find month four, replace the 4522 by the answer of 4458.79 http://www.wolframalpha.com/input/?i=4458.79%281%2B.156%2F12%29-122

OpenStudy (anonymous):

i bet there is an easier way to do this if you have an example in a book try that method

OpenStudy (anonymous):

oh so i have to do this 12 times?

OpenStudy (anonymous):

here is month 6 http://www.wolframalpha.com/input/?i=4329.88%281%2B.156%2F12%29-122

OpenStudy (anonymous):

keep going gotta run

OpenStudy (anonymous):

ok thank you so much